OpenStudy (hannaha):

MEDAL AND FAN! A swimming pool is 20 yards by 10 yards and is 5 feet deep. How many square feet are in the surface area of the sides and the bottom of the pool? Remember, a swimming pool has no top.

OpenStudy (anonymous):

200 + 50 + 50 + 100 + 100 500 ftsq

OpenStudy (hannaha):

And please tell how you get the answer! thanks

OpenStudy (anonymous):

So just find the area of each side 10 x 20 = the bottom 5 x 10 = side 1 5 x 10 = side 2 5 x 20 = side 3 5 x 20 = side 4 add them all together

OpenStudy (hannaha):

Sorry that was wrong :/

OpenStudy (nehemias123):

it is 1000sqf

OpenStudy (nehemias123):

10*20=200*5=1000

OpenStudy (nehemias123):

do you get it

OpenStudy (anonymous):

Wait sorry sorry its asking for feet but youre giving in yards so there are three feet in a yard so just multiply everything by 3

OpenStudy (nehemias123):

i know

OpenStudy (hannaha):

Okay, @NEHEMIAS123 1,000 was wrong too.

OpenStudy (nehemias123):

how 3 if it is 5 feet deep so you x 200 by 5

OpenStudy (anonymous):

so its 1500 ft

OpenStudy (anonymous):

well its 500yds, but it wants feet, so you multiply by 3

OpenStudy (nehemias123):

ok forget it i give up

OpenStudy (nehemias123):

it is 1500 ft

OpenStudy (nehemias123):

i whant to go mager pool it magers 10*20*5deep and it whas 1500 sqf

OpenStudy (nehemias123):

by i am gowing

OpenStudy (hannaha):

I'm sorry guys, i know it's confusing but thanks for trying to help. The answer is 2700 but i have no idea how that is.

OpenStudy (ribhu):